OpenStudy (anonymous):

A convex polyhedron is formed by square sides and four triangular faces. The sum of the angles of all the faces is equal to 12 right angles. What is the number of edges of this polyhedron?

OpenStudy (anonymous):

I don't even know how to start, pls help me

OpenStudy (anonymous):

12 * 4 = 48 ???

OpenStudy (anonymous):

No. options are 8, 6, 4, 2 and 1.

OpenStudy (anonymous):

or wait 12 * 3 = 36

OpenStudy (anonymous):

o nvm then lol

OpenStudy (anonymous):

thx anyway xD

OpenStudy (anonymous):

I can help you with your other question

OpenStudy (kinggeorge):

Couldn't this be a square pyramid? A square bottom with 4 right angles, and 4 triangles meeting at a single point with 180 degrees, or two right triangles each. This means \(4+4*2=12\) right angles. Also, we would then have 8 edges. 4 on the square, and 4 more joining at the tip. So you would have 8 edges.

OpenStudy (anonymous):

oh that's right! thank you so much KingGeorge

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

o i see

OpenStudy (anonymous):



OpenStudy (anonymous):

8 edges :D